# Data-Efficient Learning with Neural Programs

### v1

@@ 4.4 RQ2: Sample Efficiency
To answer RQ2, we evaluate the sample efficiency of ISED against REINFORCE, IndeCateR, and IndeCateR+ on adding MNSIT digits. IndeCateR+ [21] is a variant of IndeCateR with a sampling method and loss computation customized for higher dimensional setting such as the addition of 16 MNIST digits. We vary the size of the input and output space (sum8, sum12, sum16) of $P$ as well as the sample count, and report the average accuracy and standard deviation obtained from 5 randomized runs (Tables 4, 11-13).
@@ Appendix A Explanation of Differences Between ISED and Baseline Methods
We explain the differences between ISED and prior techniques using the simple example of sum2, where digits are restricted to be between 0-2. Suppose that we are training a neural model $M_{\theta}$ for this task, and we are considering the symbol-output sample where the ground truth symbols are 1 and 2, i.e., $r_{1}=1,r_{2}=2$, and the ground truth output is $y=3$. Suppose that the predicted distributions from $M_{\theta}$ for $r_{1}$ and $r_{2}$ are $[0.1,0.6,0.3]$ and $[0.2,0.1,0.7]$ respectively. We now explain how different methods perform their loss computations.
@@ A.2 REINFORCE
Suppose REINFORCE is also initialized with a sample count of 3, and it samples the same symbol-output pairs as before. The final reward is computed by element-wise multiplying the log probability of each sample with its reward value and taking the mean, as follows:
$\begin{bmatrix}\log(0.6)+\log(0.7)\\ \log(0.6)+\log(0.2)\\ \log(0.3)+\log(0.1)\end{bmatrix}*\begin{bmatrix}1.0\\ 0.0\\ 1.0\end{bmatrix}$
@@ A.3 IndeCateR and NASR
The loss computation for IndeCateR and NASR are identical to that of REINFORCE, also providing weak signals. Furthermore, both set the reward to 0 for samples leading to incorrect predictions, effectively ignoring them, unlike ISED which penalizes such symbols. Since only the correct symbol contribute to the final reward, the signal is sparser than ISED, making it sample-inefficient.
@@ A.4 A-NeSI
A-NeSI trains two additional neural networks: a prediction model $Q_{\theta^{\prime}}$ as a surrogate for $P$, and a prior model $R_{\alpha}$ for learning the parameters $\alpha$ for the Dirichelet distribution $D_{\alpha}$.
@@ A.6 Scallop
Suppose Scallop is initialized to use the diff-top-1-proofs semiring. This means that for each possible output, Scallop will use the proof of that output with the highest probability. For instance, the most likely proof for $y=1$ is $r_{1}=1$ and $r_{2}=0$, and the probability of the output $y=1$ is $0.6*0.2$. The final vector calculated would be
@@ Appendix B Evaluation Setup
For tasks with the MNIST dataset as unstructured data, we employ LeNet [11], a 2-layer CNN-based model, except for sum8, sum12, and sum16 tasks where we choose a smaller 2-layer CNN used by IndeCateR [21]. For HWF, we also use a 2-layer CNN-based model. For leaf classification tasks, images are scaled down and passed to a simple CNN-based network with 4 convolutional layers. For the scene recognition, we use YOLOv8 and a 3-layer convolutional network for neural program methods, 7-layer CNN for the purely neural solution, and CLIP with ViT-B/32.
We use the Adam optimizer with the best learning rate among {1e--3, 5e--4, 1e--4}. We train for maximum 100 epochs, but stop early if the training saturates. For MNIST-R tasks, we used learning rate 1e--4 and trained ISED for 10 epochs, REINFORCE and IndeCateR for 50 epochs, and A-NeSI and NASR for 100 epochs. We trained ISED for 30 epochs, A-NeSI for 100 epochs, and the rest for 50 epochs for HWF and Leaf Classification with learning rate 1e--4. For Scene Recognition, we trained A-NeSI and the purely neural baseline 50 epochs and the rest 100 epochs with learning rate 5e--4. For tasks sum8 to sum16 we trained ISED for 50 epochs and the rest for 100 epochs with learning rate 1e--3. For Visual Sudoku, we follow the setting in NASR [5] and train for 10 epochs with learning rate 1e--5.
@@ B.1 Neural-GPT Experiment Prompts
Furthremore, Scenes = {bathroom, bedroom, dining room, living room, kitchen, lab, office, home lobby, basement} and Detected Objects is a list of maximum length 10 with duplicates.
@@ Appendix C Full Performance Summary
Appendix D License Information
For implementing baselines, we adapt code from the official repositories of DeepProblog [14] (Apache 2.0), Scallop [8] (MIT), A-NeSI [24] (MIT), NASR [5] (MIT), and IndeCateR [21] (Apache 2.0). Additionally, our benchmarks use Multi-illumination dataset [16] (CC BY 4.0), HWF dataset (CC BY-NC-SA 3.0) from NGS [12], a subset of the leaf database [4] (CC BY 4.0), YOLOv8 (AGPL-3.0) and CLIP (MIT).

### v2

@@ 4.4 RQ2: Sample Efficiency
To answer RQ2, we evaluate the sample efficiency of ISED against REINFORCE, IndeCateR, and IndeCateR+ on adding MNIST digits. IndeCateR+ [21] is a variant of IndeCateR with a sampling method and loss computation customized for higher dimensional setting such as the addition of 16 MNIST digits. We vary the size of the input and output space (sum8, sum12, sum16) of $P$ as well as the sample count, and report the average accuracy and standard deviation obtained from 5 randomized runs (Tables 4, 11-13).
@@ 7 Conclusion
8 Acknowledgements
We thank the anonymous reviewers for useful feedback. This research was supported by ARPA-H grant D24AC00253-00, NSF award CCF 2313010, and by a gift from AWS AI to ASSET (Penn Engineering Center on Trustworthy AI).
@@ Appendix A Explanation of Differences Between ISED and Baseline Methods
We explain the differences between ISED and prior techniques using the simple example of sum2, where digits are restricted to be between 0-2. Suppose that we are training a neural network $M_{\theta}$ for this task, and we are considering the symbol-output sample where the ground truth symbols are 1 and 2, i.e., $r_{1}=1,r_{2}=2$, and the ground truth output is $y=3$. Suppose that the predicted distributions from $M_{\theta}$ for $r_{1}$ and $r_{2}$ are $[0.1,0.6,0.3]$ and $[0.2,0.1,0.7]$ respectively. We now explain how different methods perform their loss computations.
@@ A.2 REINFORCE
Suppose REINFORCE is also initialized with a sample count of 3, and it samples the same symbol-output pairs. The final reward is computed by element-wise multiplication of the log probability of each sample with its reward value and taking the mean, as follows:
$\frac{1}{3}*\begin{bmatrix}\log(0.6)+\log(0.7)\\ \log(0.6)+\log(0.2)\\ \log(0.3)+\log(0.1)\end{bmatrix}*\begin{bmatrix}1.0\\ 0.0\\ 1.0\end{bmatrix}$
@@ A.3 IndeCateR and NASR
The loss computation for IndeCateR and NASR are identical to that of REINFORCE, also providing weak signals with fewer samples. Furthermore, both set the reward to 0 for samples leading to incorrect predictions, effectively ignoring them, unlike ISED which penalizes such symbols. Since only the correct symbol contribute to the final reward, the signal is sparser than ISED, making it sample-inefficient.
@@ A.4 A-NeSI
A-NeSI trains two additional neural networks: a prediction model $Q_{\theta^{\prime}}$ as a surrogate for $P$, and a prior model $R_{\alpha}$ for learning the parameters $\alpha$ for the Dirichlet distribution $D_{\alpha}$.
@@ A.6 Scallop
Suppose Scallop is configured to use the diff-top-1-proofs semiring. This means that for each possible output, Scallop will use the proof of that output with the highest probability. For instance, the most likely proof for $y=1$ is $r_{1}=1$ and $r_{2}=0$, and the probability of the output $y=1$ is $0.6*0.2$. The final vector calculated would be
@@ Appendix B Evaluation Setup
For tasks with the MNIST dataset as unstructured data, we employ LeNet [11], a 2-layer CNN-based model, except for sum8, sum12, and sum16 tasks where we choose a smaller 2-layer CNN used by IndeCateR [21]. For HWF, we also use a 2-layer CNN-based model. For leaf classification tasks, images are scaled down and passed to a simple CNN-based network with 4 convolutional layers. For the scene recognition, we use YOLOv8 and a 3-layer convolutional network for neural program methods, 7-layer CNN for the purely neural solution, and CLIP with ViT-B/32. For all tasks included in RQ1, other than sudoku, we remove the final softmax function at the end of each network when evaluating IndeCateR since its sampling procedure yields optimal results without the softmax. We also do that same with REINFORCE if it results in higher accuracy. Since sudoku uses a pretrained CNN, we use the same CNN across all methods, including IndeCateR and REINFORCE.
We use the Adam optimizer with the best learning rate among {1e--3, 5e--4, 1e--4}. We train for maximum 100 epochs, but stop early if the training saturates. For MNIST-R tasks, we used learning rate 1e--4 and trained ISED for 10 epochs, REINFORCE and IndeCateR for 50 epochs, and A-NeSI and NASR for 100 epochs. We trained ISED for 30 epochs, A-NeSI for 100 epochs, and the rest for 50 epochs for HWF and Leaf Classification with learning rate 1e--4. For the Scene Recognition task, we trained A-NeSI and the purely neural baseline 50 epochs and the rest 100 epochs with learning rate 5e--4. For tasks sum8 to sum16 we trained ISED for 50 epochs and the rest for 100 epochs with learning rate 1e--3. For Visual Sudoku, we follow the setting in NASR [5] and train for 10 epochs with learning rate 1e--5.
@@ B.1 Neural-GPT Experiment Prompts
Furthermore, Scenes = {bathroom, bedroom, dining room, living room, kitchen, lab, office, home lobby, basement} and Detected Objects is a list of maximum length 10 with duplicates.
@@ Appendix C Full Performance Summary
Appendix D License Information
For implementing the baselines, we adapted the code from the official repositories of DeepProblog [14] (Apache 2.0), Scallop [8] (MIT), A-NeSI [24] (MIT), NASR [5] (MIT), and IndeCateR [21] (Apache 2.0). Additionally, our benchmarks use Multi-illumination dataset [16] (CC BY 4.0), HWF dataset (CC BY-NC-SA 3.0) from NGS [12], a subset of the leaf database [4] (CC BY 4.0), YOLOv8 (AGPL-3.0) and CLIP (MIT).